calculate the gravitational force on a 65 kg weight on the moon. The mass of the Moon is 7.35x10^22 kg and its radius is 1.74x10^6 m. 


1
Expert's answer
2021-08-12T06:20:12-0400

The gravitational force

"F=G\frac{mM}{r^2}"

"F=6.67*10^{-11}*\frac{65*7.35*10^{22}}{(1.74*10^6)^2}=105\:\rm N"